OpenStudy (sam_pi):

Hi i want help in solving polynomial equation 2x^3-7x^2-15x=0 i know that the first part of the solution is x(2 x^2+3x+-10 x-15) but then i don't know how to complete solving the problem i want detailed explanation to how to solve the problem i also now the final answer is 0,-3/2 , 5 but i don't know how they found this solution

OpenStudy (kirbykirby):

\[2x^3-7x^2-15x=0\\ x(2x^2-7x-15)=0 \] This means that either \(x=0\) or \(2x^2-7x-15 =0\) Notice that the part inside the parentheses is a polynomial of degree 2 which we know how to solve. You can try factoring it: \( 2x^2-10x+3x-15=0\\ 2x(x-5)+3(x-5)=0\\ (x-5)(2x+3)=0\) which means: \(x-5=0\) or \(2x+3 = 0\) \(\implies x=5, \text{ or }\\ 2x=-3 \implies x=-3/2\)

OpenStudy (goformit100):

Hello, and A Warm Welcome to Open Study!

OpenStudy (sam_pi):

ok , in the beginning why did you place the -10 before the 3 ? and here 2x(x-5)+3(x-5)=0 how did you factor this did you divide the 10 /2 to get 5 and 15/3 to get 5 ? why did you do that ?

OpenStudy (kirbykirby):

I broke up \(-7x\) into the sum of \(-10x \) and \(3x\) Why do we do this? When you have a polynomial of the form \(ax^2 + bx + c\), where \(a\) is not just 1, then you find 2 numbers such that: \(m \times n = a \times c\) and \(m + n = b\) Your polynomial is \(2x^2 - 7x - 15\), so \(a=2 \\b=-7 \\c=-15\) So, you look for 2 numbers \(m, n \) such that these will hold: \(m \times n = 2 \times (-15)=-30\) \( m + n = -7\) two such numbers that work are -10 and 3, since \(-10 \times 3 = -30\) and \(-10 + 3 = - 7\), so m = -10, n = 3 Now your polynomial \(ax^2 + \color{red}{bx} + c\) becomes: \( ax^2 + \color{red}{mx+ nx}+c \) \(= 2x^2 \color{red}{-10 x + 3 x} -15 \) Now look at the first 2 terms, and the last two terms, and factor out what you can. THat's what I did in doing \(2x(x-5)+3(x-5)\) Now to factor this out, let's look at a different but similar example. If you had: \(2x\color{blue}{y} + 3\color{blue}{y} \) I hope you can see that you can factor it as \(\color{blue}{y}(2x+3\)) Now just go back to our original equation and substitute \(y\) for \(x-5\) \(2x\color{blue}{(x-5)} + 3\color{blue}{(x-5)} \) factor it out the same way as we did above" \(\color{blue}{(x-5)}(2x+3\))